OpenStudy (anonymous):

HELP!!!:Convert the fraction to a decimal 8/15 A. 0.53 B. 0.815 C. 5.3 D. 8.15

OpenStudy (anonymous):

\[\frac{ 8 }{ 15 }=\frac{ x }{ 100 }=>15x=800=>x=53.3=>\frac{ 53.3 }{ 100 }=0.53\]

OpenStudy (anonymous):

Wow thanks you are absolutely amazing!!! :) @hwtrain

OpenStudy (anonymous):

No problem, you'll earn it eventually :)

OpenStudy (anonymous):

Is there another problem?

OpenStudy (anonymous):

what mixed number is equivalent to this decimal 8.00027 A. 8 27/100,000 b. 8 10,000 c.8 27/1000 d.8 27/100 @hwtrain

OpenStudy (anonymous):

Sorry for late reply, Well what you can do is just calculate the options, eventually you will get 8.00027 it should be A. because 27/100,000=0.00027 then add 8 to it and you will get 8.00027.